# Design of a LYSO Crystal Electromagnetic Calorimeter for DarkSHINE Experiment

### v1

@@ IV.3 Trigger efficiency
Table 9 shows the trigger efficiency with the energy of different regions. By using only 125 crystals in Region-I, more than 99% of the inclusive backgrounds can be filtered out. When using the sum of the energy from all crystals in the ECAL for triggering, an efficiency of nearly 100% can be achieved, except for several rare background processes. For the dark photon signal, this trigger will not cause any impact because the actual signal region requires the total energy in the ECAL to be less than 2.5 GeV.
V Radiation damage
@@
Given the high-energy and high-frequency beam environment, the ECAL, particularly its central region, is subjected to significant radiation doses that may degrade its performance. Therefore, the radiation damage to the ECAL must be evaluated, focusing on crystal damage primarily from ionizing energy loss and silicon sensor damage from non-ionizing energy loss. Simulations were conducted using Geant4 to estimate the radiation damage to crystals and silicon sensors over one year.
Figure 19: Distribution of radiation damage in ECAL area over one year of operation (a) Radiation dose absorbed by crystals, with a maximum value of $10^{7}$ rad. (b) Equivalent 1 MeV neutron fluence in silicon sensors, with a maximum value of $10^{13}$ Neq.
@@
Radiation damage to crystals mainly results from ionizing energy loss of incident particles, evaluated by the absorbed dose—the ionizing energy absorbed per unit volume of the crystal. In the simulation, the dose absorbed by each crystal per year of detector operation was recorded. For the crystals receiving the maximum dose, the absorbed dose is about $10^{7}$ rad (Figure 19(a)). Most inorganic scintillators commonly used in high-energy physics detectors, like CsI, BGO, and PWO, lose significant light yield after such a dose. However, LYSO exhibits only a small reduction in light yieldInorganicScintillators , meeting the radiation resistance requirements for the DarkSHINE ECAL. However, ionizing energy loss within the crystal can also lead to uneven response and fluorescent effects, which need more investigation and test.
@@
Radiation damage to silicon sensors mainly results from non-ionizing energy loss, typically expressed using an equivalent 1 MeV neutron fluence. First, the cumulative non-ionizing energy loss $E_{1}$ accumulated by each sensor over one year was simulated. Then, the average non-ionizing energy deposition $E_{2}$ for 1 MeV neutrons passing through one silicon sensor was calculated. The ratio $E_{1}/E_{2}$ represents the equivalent neutron fluence for one year’s operation. In the DarkSHINE ECAL, the equivalent 1 MeV neutron fluence of silicon sensors in the most heavily irradiated area is about $10^{13}$ (Figure 19(b)). This significant dose could cause the dark current of most commercial silicon sensors to increase by several orders of magnitude, rendering them unusable. Therefore, silicon sensors with excellent radiation resistance are required, as indicated in studiesSiPM_1 ; SiPM_2 ; SiPM_3 .
VI Conclusion
@@
This paper details the design and optimization of the LYSO crystal-based electromagnetic calorimeter (ECAL) for the DarkSHINE experiment, aimed at detecting dark photons, potential candidates for dark force mediators. The study involved comprehensive simulations and evaluations to identify an optimal ECAL configuration, balancing performance, cost, and durability.
The chosen crystal size of 2.5$\times$2.5$\times$4 cm3 emerged from a meticulous analysis of energy containment and resolution for electrons with energies between 1 and 8 GeV. This crystal size ensures high energy resolution and effective energy absorption while maintaining a manageable number of channels for data acquisition. The ECAL’s overall dimensions of 52.5$\times$52.5$\times$44 cm3 were selected to provide a substantial signal efficiency while considering construction and operational costs.
The energy deposition patterns within the ECAL were studied, revealing that crystals near the center absorbed significantly more energy than those at the periphery. This understanding led to the determination of a suitable dynamic range for energy measurements, with a 4 GeV energy limit per crystal found to be adequate. This limit ensures that the ECAL can accurately measure energy deposits while filtering out the majority of background processes. Additionally, potential triggering strategies based on energy distribution were proposed to enhance data processing efficiency.
@@
The radiation tolerance of the ECAL components was also assessed, highlighting the suitability of LYSO crystals due to their small light yield reduction under high radiation doses. The silicon sensors selected for the ECAL should also have excellent resistance to radiation-induced damage, crucial for maintaining detector performance in the high-radiation environment anticipated for the DarkSHINE experiment.
In summary, the ECAL design for DarkSHINE incorporates a robust configuration of LYSO crystals and radiation-resistant silicon sensors, optimized through extensive simulation studies. This design ensures high sensitivity to dark photon signals, efficient background rejection, and reliable performance under challenging conditions. The detailed simulation and optimization efforts have culminated in a matured and robust ECAL design to greatly substantiate the future DarkSHINE experiment.

### v2

@@ V.3 Trigger efficiency
In Table 6, with a 4 GeV trigger, most of the inclusive background will be rejected, while dark photon signals, especially those with large masses, will suffer only a small loss. By using only 125 crystals in Region-I, over 99.9% of the inclusive backgrounds can be filtered out, which can effectively conserve resources such as bandwidth and storage. When using the sum of the energy from all crystals in the ECAL for triggering, an efficiency of nearly 100% can be achieved, except for several rare background processes. The trigger strategy shows promise in significantly reducing background events and a small impact on dark photon signals, even when utilizing only the channels in Region-I. It will serve as an important reference for the electronics design of future detector.
VI Energy Resolution
VI.1 Energy Digitization
Accurate simulation of the detector response requires a precise description of digitization effects. Digitization applies a series of realistic effects to the simulation results, mimicking the behavior of an actual detector and bringing the simulation more comparable to real experimental conditions. The energy directly obtained from Geant4, known as the truth energy, represents the ideal energy deposition of a particle in a perfect detection scenario, devoid of any detector effects or measurement errors. Thus, digitization provides a more realistic depiction of detector performance, offering stronger evidence for design. Zhu:2024jrf ; Li:2023
The main goal of energy digitization is to parameterize the behavior of each element related to energy measurement. This involves applying smearing based on the preliminary experiments. For the DarkSHINE ECAL, the energy digitization process can be primarily divided into three mian parts: scintillation digitization, SiPM digitization, and ADC digitization. These parts correspond to the scintillator’s light emission and decay, the SiPM response, and the behavior of the readout electronics, respectively, and are based on preliminary experimental measurements.
VI.1.1 Scintillation digitization
The first step is scintillation digitization. In this part, we sampled the fluctuations that may occur during the generation and attenuation of scintillation light. We set an intrinsic light yield of 30,000 photons per MeV for the LYSO crystal scintillator, assuming a 10% fluctuation across all scintillators. The scintillation light is proportional to the energy deposited in the crystal and assumed to follow a Poisson distribution. The scintillation light experiences attenuation during propagation within the crystal and detection by the SiPM. The attenuation depends on the photon transportation length and photon detection efficiency (PDE) of the SiPM. The measured light yield helps to determine the amount of light loss during propagation and detection. Although the measured light yield for all crystal-SiPM units shows fluctuations, these can be calibrated to some extent, assumed here to have a 1% accuracy. The parameters used in this step are shown in Table 7.
Table 7: Parameters in scintillation digitization.
Scintillator Light yield (intrinsic) Light yield fluctuation
LYSO 30000 ph/MeV 10%
Light yield calibration accuracy Light yield (measured)
1% 150 p.e./MeV
VI.1.2 SiPM digitization
For SiPM digitization, a toy Monte Carlo model zhao2024 simulates the SiPM’s response to scintillation light from a LYSO crystal. This model is specifically designed to describe the relationship between the number of photons detected by the SiPM and the number of incident photons. The simulation is based on the HAMAMATSU S14160-3010PS datasheet S14160-3010PS , featuring a 3×3 mm2 sensitive area, 10 $\mu$m pixels, and nearly 90,000 pixels in total, with a LYSO crystal measuring 2.5$\times$2.5$\times$4 cm3. The model accounts for the PDE, pixel density, pixel recovery, and crosstalk effects of the SiPM, as well as the scintillation decay time, transmittance, emission spectrum, and absorption spectrum of the LYSO crystal. The parameters used in the SiPM digitization are listed in Table 8.
Table 8:  Parameters in SiPM digitization.
Active area (mm2) Pixel pitch ($\mu$m) Pixel number
3.0$\times$3.0 10 89984
PDE ($\lambda=\lambda_{p}$) Fill factor Gain fluctuation
18% 31% 10%
As shown in Figure 13, $N_{fired}$ represents the number of photons detected by the SiPM, $\epsilon$ denotes the PDE of SiPM, and $N_{In}$ refers to the number of photons incident on the SiPM. The SiPM exhibits a nearly linear response when the effective photon count, calculated as the product of the PDE ($\epsilon$) and the number of incident photons ($N_{In}$), is below 10,000. As the effective photon counts increases, the response gradually deviates from linearity and approaches saturation. A formula from kotera2016describing is employed to fit the SiPM’s response, which aids in modeling its behavior for varying numbers of incident photons, and provides a means to correct for saturation effects.
In the digitization of SiPM, the number of photoelectrons in the scintillation digitization is used as the input for the fitting function shown in Figure 13, resulting in a response that includes the non-linear effects of the SiPM. The fluctuations during this process can be also obtained, which is approximately proportional to the square root of the effective photon counts. Furthermore, the inverse of this fitting function can be applied to correct the non-linear response of SiPM, while this correction will inevitably introduce fluctuations once again. PIEMONTE20192 ; NAGY201444
Figure 13:  The simulated response of the SiPM to scintillation light from a 2.5$\times$2.5$\times$4 cm3 LYSO crystal is shown. This SiPM-LYSO setup is identical to that used in the DarkSHINE ECAL. Performance parameters of the SiPM, including PDE, fill factor, pixel pitch, pixel counts, and recovery time, are referenced from the HAMAMATSU S14160-3010PS datasheet S14160-3010PS . The detected photons exhibit a time structure influenced by the scintillation decay time of LYSO and the geometric effects of the crystal, which is considered in the simulation. The horizontal axis represents the effective photon counts, calculated as the product of the PDE ($\epsilon$) and the number of incident photons ($N_{In}$). The upper figure displays the variation in the number of photons detected by the SiPM ($N_{fired}$) as a function of the effective photon counts. The lower figure illustrates the degree to which the SiPM’s response deviates from linearity. It can be observed that for effective photon counts below 10,000, the SiPM’s response remains largely linear. However, as the effective photon counts exceeds 10,000, the response begins to deviate from linearity and gradually saturates.
VI.1.3 ADC digitization
The behavior of the readout electronics is simulated in the ADC digitization process. Table 9 shows the parameters used in ADC digitization. Here, we assume that the charge produced by the SiPM is recorded by a 12-bit multi-channel analyzer with a total of 4096 ADC counts. The gain is defined as the number of ADC counts output per single photoelectron signal. There are three gain modes to expand the dynamic range of energy measurement: high gain, medium gain, and low gain. The energy dynamic range for high gain is approximately 0 to 2.7 MeV, primarily designed for energy calibration using radioactive sources. The maximum energy measured by the designed electronics is approximately 5461 MeV, which is higher than the 4 GeV result shown in Section V. This increase is due to a series of smearing effects during digitization, which can result in energies greater than the original values. In high gain and medium gain modes, when the ADC exceeds 4000, it will automatically switch to the next gain level. A 3-ADC DAQ noise was set for each gain mode, representing the intrinsic noise of the data acquisition system. The SiPM noise is set to 1 ADC at high gain mode, and varies with gain. The digitization model characterizes the SiPM noise as the average dark noise present in each event’s waveform. This accounts for the scintillation decay time of LYSO and the dark count rate of the SiPM. Additionally, a 1% calibration accuracy was also assumed during photoelectron measurement.
Table 9: Parameters in ADC digitization.
Modes Gain ADC/p.e. Charge calibration accuracy
High gain $\times$2000 10 1%
Medium gain $\times$40 0.2 1%
Low gain $\times$1 0.005 1%
SiPM noise DAQ noise Switching point Energy range
1 ADC 3 ADC 4000 ADC 0-2.7 MeV
0.02 ADC 3 ADC 4000 ADC 2.7-133.3 MeV
0.0005 ADC 3 ADC - 133.3-5461 MeV
VI.2 Single channel performance
After digitization, the ideal energy deposited in the crystal within Geant4 is transformed into a signal that more closely represents the actual response of the detector. The Equivalent Noise Energy (ENE) for the electronics of a single channel digitization is shown in Figure 14(a). The ENE is calculated using the following formula:
$ENE=\frac{Noise}{ADC_{p.e.}\times LY_{mea}}$ (1)
where $Noise$ is the number of ADC counts of noise, representing the sum of SiPM noise and DAQ noise, as listed in Table 9. $ADC_{p.e.}$ is the number of ADC counts per photoelectron, also provided in Table 9. $LY_{mea}$ is the number of photoelectrons detected by the SiPM coupled to the crystal when the crystal absorbs 1 MeV of energy. In our setup, $LY_{mea}$ is set to 150 p.e./MeV (Table 7).
In Figure 14(a), three distinct ENE regions correspond to different gain modes. Since ENE represents the electronics noise and is independent of the input signal, it remains constant across the energy range covered by each electronics gain. However, at the transition points between these ranges, the ENE exhibits jumps, as the gain changes as well as $Noise$ and $ADC_{p.e.}$.
Figure 14:  (a) Equivalent noise energy (ENE) of a single channel. The x-axis represents the input energy in digitization model, while the y-axis represents ENE. The three bins in the figure correspond to the three gain modes, each with a fixed ENE value that is independent of input energy. The width of each bin represents the dynamic range of the energy measurement at that gain. (b) Energy resolution of a single channel. The x-axis represents the input energy in digitization model, and y-axis indicates the channel’s energy resolution, calculated as the ratio of the sigma to the mean value from a Gaussian fit of the output energy distribution. At energies below 100 MeV, the primary contributor to resolution is the SiPM digitization, since it depends on efficiencies of light collection and photon detection, which introduced larger fluctuations in monte carlo sampling than intrinsic light output and ADC noise. As energy increases, the scintillation digitization and ADC digitization become more dominant, mainly due to the constant terms reflecting calibration accuracy.
After incorporating the three digitization effects, the single-channel energy resolution is shown in Figure 14(b). We individually simulated the digitization process for each component to obtain its resolution and also simulated the process with all three components present to determine the overall resolution. The scintillation digitization consists of a Poisson sampling based on the intrinsic light yield of LYSO crystal, along with a series of Gaussian samplings. As a result, this part of the energy resolution decreases with increasing incident energy. But, due to the high intrinsic light yield of the crystal, this change is not very noticeable in Figure 14(b) and is instead dominated by the constant term introduced by the Gaussian sampling, which reflects the calibration accuracy. The resolution of SiPM digitization also decreases with increasing energy. However, since its statistical sampling is based on the number of detected photoelectrons, which is significantly lower than the number of generated scintillation photons due to attenuation and the PDE of SiPM, the resolution of SiPM digitization is worse than that of scintillation digitization. Moreover, applying a non-linearity correction further degrades the resolution, as the correction process introduces additional fluctuations. For ADC digitization, the energy resolution shows two jumps at 5 MeV and 500 MeV, caused by an increase in ENE due to gain switching. Additionally, a constant term appears in the high-energy region, reflecting the calibration accuracy assumed in the model. Overall, at lower energies (less than 100 MeV), the resolution of the SiPM digitization is the main contributor. At higher energies, contributions from the scintillation digitization and ADC digitization become more significant.
The energy resolutions in Figure 14(b) were obtained based on parameters in Table 7, 8 and 9, which referred to the preliminary experimental measurements. Optimizing the LYSO-SiPM units, such as improving light collection efficiency to get a larger light output, can enhance the channel resolution.
VI.3 ECAL performance
The performance of the full detector can be determined by applying digitization to each channel. Figure 15 show the energy resolution and energy containment of the ECAL. $E_{Truth}$ is the energy of incident electrons. In simulation, the ECAL is configured to consist of 21$\times$21$\times$11 LYSO crystals, each with a volume of 2.5$\times$2.5$\times$4 cm3, consistent with the results from Section IV. The target and other detectors are not included. The incident particles are electrons with energies ranging from 1 to 8 GeV, hitting the center of ECAL’s front face within a circular area with a radius of 3 cm.
Figure 15:  Energy resolution and energy containment of the ECAL for 1–8 GeV incident electrons, without the target and other detectors. The x-axis represents the energy of the incident electrons. The top figure displays the ECAL’s energy resolution, with the fitted stochastic term achieving better than 2%. The bottom figure shows the energy containment, defined as the ratio of the energy deposited in the ECAL to the incident energy.
The energy resolution is plotted at the top of Figure 15. After digitization, the statistical term of ECAL’s energy resolution remains better than 2%, thanks to the very high intrinsic light yield of LYSO crystal. The high energy resolution will allow for more precise measurement of recoil electron energy and enhance the reliability in distinguishing between signal and background. The overall energy resolution of the ECAL is typically better than that of a single channel (Figure 14(b)) due to statistical averaging, where uncertainties from individual channels tend to cancel out, and energy sharing across multiple channels, which reduces the impact of measurement errors. Additionally, variations in the response of different channels are compensated when combined, improving the overall resolution.
The bottom of Figure 15 shows the energy containment of ECAL, which is defined as the ratio of deposited energy in ECAL to the incident energy. The energy containment is less than 100%, mainly because a small portion of the energy is deposited in passive materials outside the crystals, such as the reflective films, PCB, and support structures, which can not be detected.
Since the performance of the ECAL is influenced by the digitization model, which is based on preliminary experimental results from a specific setup, optimizing the detector configuration could change the performance. However, the key objective is to develop a tool that can more accurately evaluate the detector’s performance, making it closer to that of a real detector.
VII Radiation damage
@@
Given the high-energy and high-frequency beam environment, the ECAL, particularly its central region, is subjected to significant radiation dose that may degrade its performance. Therefore, the radiation damage to the ECAL must be evaluated, focusing on crystal damage primarily from ionizing energy loss and silicon sensor damage from non-ionizing energy loss. Simulations were conducted using Geant4 to estimate the radiation damage to crystals and silicon sensors under $3\times 10^{14}$ electrons-on-target events, corresponding to one year of operation at a 10 MHz repetition rate.
Figure 16:  Distribution of radiation damage in the ECAL region under $3\times 10^{14}$ electrons-on-target events, corresponding to one year of operation at a 10 MHz repetition rate. The two distributions illustrate the radiation damage along the ECAL symmetry plane in beam direction. In figures, the horizontal axis represents the beam direction, while the vertical axis corresponds to the transverse directions. Each segment indicates the position of a crystal. (a) Total Ionizing Dose (TID) absorbed by crystals, with a maximum value of $10^{7}$ rad. (b) Non-Ionizing Energy Loss (NIEL) in silicon sensors, expressed as the equivalent 1 MeV neutron flux, with a maximum value of $10^{13}$ per square centimeter.
@@
Radiation damage to crystals mainly results from ionizing energy loss of incident particles, which can be evaluated by the Total Ionizing Dose (TID). The TID in crystal is defined as the ionizing energy absorbed per unit mass of crystal. The average ionizing energy loss absorbed by each crystal in the ECAL per event was calculated based on one million electrons-on-target events. This average value was then multiplied by $3\times 10^{14}$ to estimate the TID for each crystal after one year of operation. For the crystal absorb the maximum dose, the value of TID is about $10^{7}$ rad (Figure 16(a)). Most inorganic scintillators commonly used in high-energy physics detectors, like CsI, BGO, and PWO, lose significant light yield after such a dose. However, LYSO exhibits only a small reduction in light yieldZhu:2019ihr , meeting the radiation resistance requirements for the DarkSHINE ECAL.
@@
Radiation damage to silicon sensors mainly results from Non-Ionizing Energy Loss (NIEL), typically expressed using the equivalent 1 MeV neutron flux. To obtain the flux on each sensor, the average NIEL per event for each sensor was first calculated through simulation, referred to as $E_{1}$. Then, the average NIEL for a 1 MeV neutron passing through a single sensor was simulated, referred to as $E_{2}$. Finally, the ratio $E_{1}/E_{2}$ provides the equivalent 1 MeV neutron flux for each sensor. In the DarkSHINE ECAL, the equivalent 1 MeV neutron flux on silicon sensors in the most heavily irradiated area is about $10^{13}$ per square centimeter (Figure 16(b)). This significant radiation could cause the dark current of general sensors to increase by several orders of magnitude, rendering them unusable. Therefore, silicon sensors with excellent radiation resistance are required for our experiment.ULYANOV2020164203 ; SANCHEZMAJOS2009506 ; PREGHENELLA2023167661
VIII Conclusion
@@
In this study, we presented the design and optimization of a LYSO crystal ECAL for the DarkSHINE experiment, aimed at detecting dark photons, potential candidates for dark force mediators. Through comprehensive simulations, we optimized the ECAL’s structure and evaluated its key performance metrics.
The final ECAL design consists of 21$\times$21$\times$11 LYSO crystals, each measuring 2.5$\times$2.5$\times$4 cm3, arranged in a staggered configuration to enhance detection efficiency. Both the transverse and longitudinal dimensions of ECAL were optimized to achieve high signal collection while maintaining cost efficiency. A 4 GeV energy dynamic range was established for each channel to ensure accurate energy measurements without saturation, which is crucial for background rejection and precise signal detection.
A dedicated digitization model parameterizing the scintillation, SiPM, and ADC behaviors was developed to provide a realistic representation of the detector’s performance. This model allowed us to study the energy resolution and containment of the ECAL, demonstrating that the statistical term of the energy resolution can be better than 2%, thanks to the high intrinsic light yield of the LYSO crystals.
@@
Radiation damage was also assessed. The TID absorbed by the crystals and the non-ionizing energy loss in the silicon sensors were estimated for one year of operation at a 10 MHz repetition rate. Due to the small light yield reduction under high radiation doses, LYSO crystal is an ideal scintillator material for the DarkSHINE experiment. The selected silicon sensors for the ECAL should also have excellent resistance to radiation-induced damage, which is critical for maintaining detector performance in the high-radiation environment anticipated for the experiment.